Costco Wholesale Corporation operates membership warehouses selling food, appliances, consumer electronics, apparel and other household goods at 721locations across the U.S. as well as in Canada, the United Kingdom, Japan, Australia, South Korea, Taiwan, Mexico and Puerto Rico. As of its fiscal year-end 2016, Costco had approximately 86.7 million members. Selected fiscal-year information from the company's balance sheets follows.

Selected Balance Sheet Data ($ millions)

2016 2015
Merchandise inventories $8,969 $8,908
Deferred membership income (liability) 1,362 1,269


(a) During fiscal 2016, Costco collected $2,739 million cash for membership fees. Use the financial statement effects template to record the cash collected for membership fees.
(b) Costco recorded merchandise costs (that is, cost of goods sold) of $102,901 million in 2016. Record this transaction in the financial statement effects template.
(c) Determine the value of merchandise that Costco purchased during 2016. Use the financial statement effects template to record these merchandise purchases. Assume all of Costco's purchases are on credit.
Answer #1

ANSWER

Balance sheet ($ in million)
Transaction Cash Assets + Non Cash Assets = Liabilities + Contributed Capital + Earned Capital
(a) $         2,739 + 0 = $         2,739 + 0 + 0
(b) 0 + $            -1,02,901 = 0 + 0 + $        -1,02,901
(c) 0 + $             1,02,962 = $   1,02,962 + 0 + 0
Income Statement ($ in million)
Transaction Revenue - Expense = Net Income
(a) 0 - 0 = 0
(b) 0 - $             1,02,901 = $ -1,02,901
(c) 0 - 0 = 0
Workings:
(c) Cost of goods sold = Opening Inventory + Purchases - sales
Purchases = $102901 - $8908 + $8969
Purchases = $102962

---------------------------------------------------------------------

DEAR STUDENT,
If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************